# Complementary Time–Space Tradeoff for Self-Stabilizing Leader Election: Polynomial States Meet Sublinear Time

### v1

@@ 1.3 Our Contribution
This paper affirmatively resolves the first open problem of Burman et al. [14], independently of and nearly concurrently with Austin et al. [9]. [^4] We posted this paper on arXiv in May 2025, the same month that [9] was released on arXiv. Specifically, we prove the following theorem. (We write $\lg x$ for $\log_{2}x$ throughout this paper.)
@@
For any positive integer $\rho$ with $1\leq\rho\leq\sqrt{n}$, there exists a protocol that solves SS-RK (and hence SS-LE) within $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ expected time using $2^{2\rho\lg\rho+O(\log n)}$ states (equivalently, $2\rho\lg\rho+O(\log n)$ bits).
Choosing any $\rho=\omega(1)$ yields a sublinear-time, sub-exponential-state protocol. For example, $\rho=\lfloor\sqrt{n}\rfloor$ achieves $O(\sqrt{n}\log n)$ expected time with $2^{(\sqrt{n}/2)\lg n}\cdot\mathrm{poly}(n)$ states, while $\rho=\Theta\left(\nicefrac{{\log n}}{{\log\log n}}\right)$ achieves $O\left(n\cdot\nicefrac{{\log\log n}}{{\log n}}\right)=o(n)$ expected time with only polynomially many states (see Table 2).
The proposed protocol and that of Austin et al. [9] exhibit complementary strengths and trade-offs. The proposed protocol uses significantly fewer states than the protocol of Austin et al. requires to achieve any expected stabilization time above $\Theta(\sqrt{n}\log n)$. However, the restriction $\rho\leq\sqrt{n}$ prevents us from attaining stabilization time $o(\sqrt{n}\log n)$. Conversely, the protocol of Austin et al. can exceed this bound; in particular it can achieve an optimal $O(\log n)$-time solution with $2^{O(n^{2}\log n)}$ states.
The second example mentioned above (i.e., the case of $\rho=\Theta\left(\nicefrac{{\log n}}{{\log\log n}}\right)$) resolves the first open question of Burman et al. [14] in stronger form, upgrading the state bound from sub-exponential to polynomial:
@@
Note that the protocol of [9] does not achieve this: it requires $\rho=\omega(\log n)$ to stabilize in sublinear time, which entails $\omega(\log^{2}n)$ bits (super-polynomial states). One might argue that Austin et al. [9] only claim an $O(\nicefrac{{n}}{{\rho}}\cdot\log n)$ bound for w.h.p. stabilization time and that a refined analysis or slight modification could reduce their expected time to $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ or $O(\nicefrac{{n}}{{\rho}})$. However, the (non-self-stabilizing) ranking component employed in their SS-RK protocol, named ${\textsc{AssignRanks}}_{\rho}$ with $21\leq\rho<n/$ 2, inherently requires $\Omega(\nicefrac{{n}}{{\rho}}\cdot\log n)$ expected stabilization time, while using $2^{O(\rho\log n)}$ states.
@@
The ranking protocol ${\textsc{AssignRanks}}_{\rho}$ of Austin et al. [9] requires $\Omega(\nicefrac{{n}}{{\rho}}\cdot\log n)$ expected stabilization time.
It is worth mentioning that the proposed SS-LE protocol employs a loosely-stabilizing leader election (LS-LE) protocol as a subroutine. LS-LE [30, 34, 28] is a weakened variant of SS-LE: from any configuration, the population quickly reaches a safe configuration from which exactly one leader persists for an arbitrarily long (but finite) expected time. The time-optimal LS-LE protocol of Sudo, Eguchi, Izumi, and Masuzawa [28], parameterized by $\tau\geq 1$, reaches a safe configuration in $O(\tau\log n)$ expected time and thereafter maintains a unique leader for $\Omega(n^{\tau})$ expected time, using only $O(\tau\log n)$ states. Since its time and space complexities are negligible, we can assume a unique leader almost always exists when designing our SS-LE protocol. To the best of our knowledge, this is the first work where an LS-LE protocol aids the design of an SS-LE protocol—an intriguing approach in which a stable leader is elected with the help of an unstable leader.
2 Preliminaries
2.1 Term and Notation
Throughout this paper, we write $[x,y]=\{i\in\mathbb{N}:x\leq i\leq y\}$ for any real numbers $x$ and $y$. We write $\lg x$ for $\log_{2}x$. For any set of configurations $\mathcal{C}$, we say the population enters $\mathcal{C}$ when it reaches some configuration in $\mathcal{C}$.
Although SS-RK was informally introduced in Section 1, we now give its formal definition.
Definition 1 (Self-Stabilizing Ranking (SS-RK)).
A protocol $\mathcal{P}$ solves SS-RK if, with probability 1, the population under $\mathcal{P}$ eventually reaches an output-stable configuration in which all agents output distinct integers in $[1,n]$.
@@ 2.1 Term and Notation
Multiple definitions of “with high probability” (w.h.p.) appear in the population-protocol literature (e.g., with probability $1-n^{-1}$ or $1-n^{-\Omega(1)}$). As in [10, 8], we adopt the following definition.
Definition 2 (with high probability).
A property of a protocol $\mathcal{P}$ holds with high probability if, for any constant $\eta=\Theta(1)$, we can ensure the property is satisfied with a probability of $1-O(n^{-\eta})$ by adjusting the constant parameters of $\mathcal{P}$, if necessary.
For example, if a protocol $\mathcal{P}_{\tau}$ with parameter $\tau=\Theta(1)$ elects a leader in $O(\tau\,n\log n)$ interactions with probability $1-n^{-c\tau}$ (for some constant $c=\Theta(1)$), we say that $\mathcal{P}_{\tau}$ elects a leader in $O(n\log n)$ interactions w.h.p. (or more precisely $O(\tau\,n\log n)$ interactions w.h.p.).
3 Self-stabilizing Ranking
@@
This section proves our main theorem (Theorem 1) by presenting the SS-RK protocol $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$. Given any integer $\rho\in[1,\sqrt{n}]$, this protocol solves SS-RK within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions (i.e., $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ time) in expectation using $2^{2\rho\lg\rho+O(\log n)}$ states.
3.1 Protocol Overview
@@
The protocol $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ comprises three sub-protocols: FindTarget, ${\textsc{Detect}}(r,\rho)$, and Rank (parametrized by $2\leq\chi\leq n$). These components serve the following roles:
@@
  • FindTarget: if some rank $r\in[1,n]$ is held by two or more agents, selects one such $r$ as the target rank with probability $1-o(1)$.
@@
  • ${\textsc{Detect}}(r,\rho)$: if the target rank $r$ is held by two or more agents, detects the collision with probability $\Omega(1)$.
@@
  • Rank: upon collision detection, reassigns unique ranks in $[1,n]$ to all agents with probability $1-o(1)$.
@@
We execute these components sequentially in cycles. To ensure the above success probabilities, the sub-protocols FindTarget, ${\textsc{Detect}}(r,\rho)$, and Rank require $O(n^{3/2}\log n)$, $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$, and $O(n^{3/2}\log n)$ interactions, respectively; the second bound dominates for all $\rho\in[1,\sqrt{n}]$. By the union bound, each cycle succeeds with probability $1-(o(1)+(1-\Omega(1))+o(1))=\Omega(1)$. Therefore, the expected number of cycles to success is $O(1)$, yielding $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ expected interactions overall.
@@
The core novelty of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ lies in ${\textsc{Detect}}(r,\rho)$. As noted in Section 1.3, repeating a collision-detection protocol with periodic resets carries a small but nonzero probability of reset failure, leaving outdated information in some agents. Such residual data causes false-positive detections and triggers rank reassignment, preventing stabilization. Consequently, the polynomial-time, polynomial-state protocol of Araya and Sudo [8] does not immediately yield a polynomial-time, polynomial-state SS-RK protocol. Austin et al. [9] resolving this issue by avoiding periodical resets: they reinitialize their collision-detection module only when a collision is detected. In contrast, our component ${\textsc{Detect}}(r,\rho)$ tolerates outdated information: once each agent has initialized its variables at least once, ${\textsc{Detect}}(r,\rho)$ incurs no false positives, even if some agents retain outdated information from cycles long past. Therefore, once a single complete cycle of the three components successfully assigns unique ranks, ${\textsc{Detect}}(r,\rho)$ produces no false positives in subsequent cycles, ensuring stabilization.

### v2

@@ 1.3 Our Contribution
This paper affirmatively resolves the first open problem of [BCC+21a], independently of and nearly concurrently with [ABF+25]. [^4] The first version of this paper was posted on arXiv in May 2025, the same month that [ABF+25] appeared on arXiv. Specifically, we prove the following theorem. (We write $\lg x$ for $\log_{2}x$ throughout this paper.)
@@
For any positive integer $\rho$ with $2\leq\rho\leq\sqrt{n}$, there exists a protocol that solves SS-RK (and hence SS-LE) in $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ expected time using $2^{2\rho\lg^{2}\rho+O(\log n)}$ states (equivalently, $2\rho\lg^{2}\rho+O(\log n)$ bits).
Choosing any $\rho=\omega(1)$ yields a sublinear-time, sub-exponential-state protocol. For example, $\rho=\lfloor\sqrt{n}\rfloor$ achieves $O(\sqrt{n}\log n)$ expected time with $2^{\sqrt{n}\lg^{2}n}\cdot\operatorname{poly}(n)$ states, while $\rho=\Theta\left(\nicefrac{{\log n}}{{\log^{2}\log n}}\right)$ achieves $O\left(n\cdot\nicefrac{{\log^{3}\log n}}{{\log n}}\right)=o(n)$ expected time with only $\operatorname{poly}(n)$ states (see Table 2).
The proposed protocol and that of [ABF+25] exhibit complementary strengths and trade-offs. The proposed protocol uses significantly fewer states than the protocol of [ABF+25] requires to achieve any expected stabilization time above $\Theta(\sqrt{n}\log n)$. However, the restriction $\rho\leq\sqrt{n}$ prevents us from attaining stabilization time $o(\sqrt{n}\log n)$. Conversely, the protocol of [ABF+25] can exceed this bound with super-exponentially many states; in particular it can achieve an optimal $O(\log n)$-time solution with $2^{O(n^{2}\log n)}$ states.
The second example mentioned above (i.e., $\rho=\Theta\left(\nicefrac{{\log n}}{{\log^{2}\log n}}\right)$) resolves the first open question of [BCC+21a] in stronger form, upgrading the state bound from sub-exponential to polynomial:
@@
Note that the protocol of [ABF+25] does not achieve this: to stabilize in sublinear time it requires $\rho=\omega(\log n)$, which entails $\omega(\log^{2}n)$ bits (i.e., super-polynomial states). One might observe that [ABF+25] only claim an $O(\nicefrac{{n}}{{\rho}}\cdot\log n)$ bound for w.h.p. stabilization time and hope that a refined analysis or slight modification could reduce their expected time to $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ or even $O(\nicefrac{{n}}{{\rho}})$. However, their SS-RK protocol employs a non-self-stabilizing ranking subprotocol, ${\textsc{AssignRanks}}_{\rho}$ (for $1\leq\rho<n/2$), which inherently requires $\Omega(\nicefrac{{n}}{{\rho}}\cdot\log n)$ expected time to assign distinct ranks to all $n$ agents while using only $2^{O(\rho\log n)}$ states.
@@
The ranking protocol ${\textsc{AssignRanks}}_{\rho}$ of [ABF+25] requires $\Omega(\nicefrac{{n}}{{\rho}}\cdot\log n)$ expected time to assign distinct ranks to all $n$ agents.
It is worth mentioning that the proposed SS-LE protocol employs a loosely-stabilizing leader election (LS-LE) protocol as a subroutine. LS-LE [SNY+12, SOK+20, SEIM21] is a weaker variant of SS-LE: from any configuration, the population quickly reaches a safe configuration from which exactly one leader persists for an arbitrarily long (but finite) expected time. The time-optimal LS-LE protocol of Sudo, Eguchi, Izumi, and Masuzawa [SEIM21], parameterized by $\tau\geq 1$, reaches a safe configuration in $O(\tau\log n)$ expected time and thereafter maintains a unique leader for $\Omega(n^{\tau})$ expected time, using only $O(\tau\log n)$ states. Since its time and space complexities are negligible, we can assume a unique leader almost always exists when designing our SS-LE protocol. To the best of our knowledge, this is the first work where an LS-LE protocol aids the design of an SS-LE protocol—an intriguing approach in which a stable leader is elected with the help of an unstable leader.
Most modern population protocols employ the epidemic protocol [AAE08] to disseminate information from a single agent to the entire population, whereas our protocol uses a slightly slower propagation mechanism to avoid false-positive collision detection. To analyze its stabilization time, we introduce a new variant of the epidemic protocol, called the labeled epidemic. While the original epidemic spreads information to all agents within $O(\log n)$ time w.h.p. and in expectation, we show that this variant spreads information to a constant fraction of the population (i.e., $\Theta(n)$ agents) in $O(\log^{2}n)$ expected time, which may be of independent interest in the study of population protocols.
1.4 Notation
Throughout this paper, we write $x,y]=\{i\in\mathbb{Z}:x\leq i\leq y\}$ for any real numbers $x$ and $y$. For any set of configurations $\mathcal{C}$, we say the population enters $\mathcal{C}$ when it reaches some configuration in $\mathcal{C}$.
@@ 1.4 Notation
2 Protocol Overview
@@
We prove our main theorem (Theorem 1) by presenting the SS-RK protocol $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$. Given any integer $\rho\in[2,\sqrt{n}]$, this protocol solves SS-RK within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions (i.e., $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ time) in expectation using $2^{2\rho\lg^{2}\rho+O(\log n)}$ states. The pseudocode for SS-RK is presented in Algorithms 1, 2, 3, and 4 on pages 1, 2, 3, and 4, respectively.
@@
The protocol $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ comprises three sub-protocols: FindTarget, ${\textsc{Detect}}(r,\rho)$, and Rank. These components serve the following roles:
@@
FindTarget if some rank $r\in[1,n]$ is held by two or more agents, selects one such $r$ as the target rank with probability 1o(1)1-o(1).
@@
${\textsc{Detect}}(r,\rho)$ if the target rank $r$ is held by two or more agents, detects the collision with probability $\Omega(1)$.
@@
Rank upon collision detection, reassigns unique ranks in $[1,n]$ to all agents w.h.p.
@@
We execute these components sequentially in cycles. To ensure the above success probabilities, the sub-protocols FindTarget, ${\textsc{Detect}}(r,\rho)$, and Rank require $O(n^{3/2}\log n)$, $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$, and $O(n^{3/2}\log n)$ interactions, respectively; the second bound dominates for all $\rho\in[2,\sqrt{n}]$. By the union bound, each cycle succeeds with probability $1-(o(1)+(1-\Omega(1))+o(1))=\Omega(1)$. Therefore, the expected number of cycles to success is $O(1)$, yielding $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ expected interactions overall.
@@
The core novelty of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ lies in ${\textsc{Detect}}(r,\rho)$. As noted in Section 1.3, repeating a collision-detection protocol with periodic resets carries a small but nonzero probability of reset failure, leaving outdated information in some agents. Such residual data causes false-positive detections and triggers rank reassignment, preventing stabilization. Consequently, the polynomial-time, polynomial-state protocol of [AS25] does not immediately yield a polynomial-time, polynomial-state SS-RK protocol. The authors of [ABF+25] resolve this issue by avoiding periodic resets: they reinitialize their collision-detection module only when a collision is detected. In contrast, our component ${\textsc{Detect}}(r,\rho)$ tolerates outdated information: once all agents have initialized their variables at least once, ${\textsc{Detect}}(r,\rho)$ incurs no false positives, even if some agents retain outdated information from cycles long past. Therefore, once a single complete cycle of the three components successfully assigns unique ranks, ${\textsc{Detect}}(r,\rho)$ produces no false positives in subsequent cycles, ensuring stabilization.